# On Minimal Achievable Quotas in Multiwinner Voting

### v1

@@ 6 Restricted Domains
Since it is computationally hard to compute $\alpha^{*}_{\Phi}(I)$ for general instances $I$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$, we look into some restricted domains in which we can efficiently find the optimal $\alpha$-value. These domains can capture natural situations where preferences are aligned along a single dimension, such as voters and candidates positioned on a political spectrum, geographic proximity in local elections, or a preference for candidates differing along a skill or topic dimension.
@@ 6.1 Party-List Profiles
A commonly studied restricted domain is that of party-list profiles, where voters’ preferences depend solely on the party affiliation of candidates.
@@
Computing $\alpha^{*}_{\text{EJR}}(I)$ for party-list instances can be done efficiently.
Proof.
$\arg\max_{p\in P}\frac{s_{p}}{w_{p,i}+1}$
where $w_{p,i}$ is the number of candidates approved by party $p$ in the committee $W$ at timestep $i$. The intuition behind the algorithm is the following: Check at every timestep which party causes the largest violation.
For the party $p$ not to get $(\alpha,w_{p,i}+1)$-cohesive, it has to hold that $\frac{n}{k}\cdot\alpha\cdot(w_{p,i}+1)>s_{p}$. Solving for $\alpha$, we get that $\alpha>\frac{s_{p}}{w_{p,i}+1}\cdot\frac{n}{k}$ where $\frac{n}{k}$ is a constant factor for all parties and can thus be ignored.
This contradicts our assumption that $\alpha^{\prime}<\alpha$. ∎
Remark 4.
$\alpha$-EJR and $\alpha$-EJR+ are equivalent for party-list instances since any EJR+ violation $(N^{\prime},c,\ell)$ corresponds to an EJR violation $(N^{\prime},A_{v},\ell)$ for any $v\in N^{\prime}$ since by definition $\lvert A_{v}\rvert\geq\ell$.
@@ 6.2 Voter-Interval Domain
It has been shown by Aziz et al. (2017) that checking whether $W$ provides EJR for an instance $I$ is coNP-complete. However, to the best of our knowledge, there is no result on efficiently finding violations for $\ell$-cohesive groups when we only consider restricted approval profiles. More specifically, we show the following lemma, which lets us efficiently check for an $\alpha$-EJR violation given an instance $I$ and committee $W$.
@@
In the VI domain, verifying whether a committee $W$ satisfies $\alpha$-EJR can be done efficiently.
Also, for JR, computationally hard problems become tractable for the VI domain, as the following theorem proves.
@@
Computing $\alpha^{*}_{\text{EJR}}(I)$ in the VI domain can be done efficiently.
Proof Sketch.
We use the greedy algorithm in Algorithm˜1, which constructs a cardinality-minimal subset of candidates that satisfies $\alpha$-JR for a given $I$ and $\alpha$. See Lemma˜B.2 for a proof.
By checking if the size of the returned committee is at most $k$, we can efficiently verify if such a committee exists for a given $\alpha$. Due to Lemma˜4.11, the number of distinct $\alpha$-values we need to consider is upper-bounded by $n$. Thus, we can efficiently construct a committee that satisfies the optimal $\alpha$-value for $I$. ∎
In Algorithm˜1, we use the observation that in the VI domain, every candidate $c$ has a leftmost and rightmost voter according to the VI order who approves them. We denote these voters by $l_{c}$ and $r_{c}$, respectively. For $i\in[n]$, the algorithm considers the set $\{v_{1},\dots v_{i}\}$ and fixes an $\alpha$-JR violation, if existent, with a "right-most" candidate according to the VI order. Intuitively, this ensures that the violation to the left of $v_{i}$ is resolved, while covering the maximum number of voters to the right, thereby postponing the next potential violation as long as possible.
@@
   $W\leftarrow W\cup\{c\}$ with $c\in A_{i}$ and
     $r_{c^{\prime}}\sqsubseteq r_{c}$ for all $c^{\prime}\in A_{i}$.
@@ 6.3 Candidate-Interval Domain
In the CI domain, verifying whether a committee $W$ satisfies $\alpha$-EJR can be done efficiently.
@@
Computing $\alpha^{*}_{\text{EJR}}(I)$ in the CI domain can be done efficiently.
@@ 7 Empirical Evaluation
Figure 5: Distribution of optimal $\alpha$-values, $\alpha_{\text{EJR}}^{*}$, under the Impartial Culture and the Euclidean Threshold model, based on 6400 generated instances each.
Figure 6: Cumulative distribution of the optimal $\alpha$-value, $\alpha_{\text{EJR}}^{*}$, and the $\alpha$-value achieved by MES, seq-Phragmén, $C$, and PAV under the Impartial Culture and Euclidean Threshold model, based on 6400 samples.
Finally, we complement our theoretical results with an empirical analysis. We generate synthetic voting instances to evaluate the resulting values of $\alpha$ and to assess the performance of existing multi-winner voting rules. Our primary focus lies on $\alpha$-EJR; additional plots for $\alpha$-JR are deferred to Appendix˜C. Figure˜5 reports the distribution of optimal values, $\alpha^{*}_{\text{EJR}}$, under the Impartial Culture (IC) and the Euclidean Threshold (EUT) model, two commonly used probabilistic models for generating approval profiles. For each model, we consider instances of varying sizes—ranging from small to moderate—with up to $60$ voters and $15$ candidates. Both models are well used in approval-based multiwinner voting (Bredereck et al., 2019; Faliszewski et al., 2017; Elkind et al., 2017). Further details on the parameter configurations of them are provided in Appendix˜C.
@@
In addition to our theoretical analysis in Section˜4.2, we evaluate the performance of existing voting rules on randomly generated instances. Figure˜6 demonstrates that MES, seq-Phragmén, and PAV typically achieve committees with comparatively low $\alpha$-values with respect to EJR, indicating that these rules perform well on average. Note that MES is completed with seq-Phragmén. Our experiments, however, also uncover instances in which all three rules yield committees whose $\alpha$-values exceed the optimum by more than a factor of four, highlighting the presence of substantial worst-case deviations even in average-case data. These cases can be mostly seen among larger instances. In Appendix˜C, we provide more fine-grained information on how the voting rules perform under different instance sizes. CC does not perform very well, which is unsurprising because its only goal is to satisfy as many voters as possible and because it does not satisfy EJR for all instances. In Figure˜8 we show that, on the contrary, CC performs best out of the considered rules for $\alpha_{JR}^{*}$.
@@ 8 Conclusion
We demonstrate that our dynamic extensions of JR and EJR can lead to much fairer outcomes for various instances. Our axioms possess a clear fairness appeal, as they seek to minimize the underrepresentation of large aligned groups, which in real-world political settings may otherwise foster resentment. However, as for many other voting rules that try to maximize some objective, like CC or PAV, computing the optimal $\alpha$-value is NP-hard, even for JR.
While existing voting rules can exhibit poor worst-case behavior theoretically, they perform reasonably well on our artificially generated datasets. A natural direction for future work is to evaluate these rules on real-world data, for instance, using the Polkadot dataset (Boehmer et al., 2024). When restricting the set of voting instances, we obtain several positive results regarding the efficient computation and verification of $\alpha$-values. Empirical results underline the importance of considering adaptive quotas in multiwinner approval voting and perhaps other settings. Further extensions could include generalizing our framework to weakly cohesive groups or developing broader notions of cohesiveness based on justified representation functions. Such generalizations may also prove useful in continuous settings, such as budget aggregation.
@@ Appendix A Mentioned Voting Rules
Selects all committees that maximize the CC-score, i.e. $\arg\max_{W\subseteq C,\lvert W\rvert=k}\text{CC}(W)$. The CC-score is equal to the number of voters approving at least one candidate in the committee. Formally,
$\text{CC}(W)=\lvert\{i\in N:A_{i}\cap W\neq\emptyset\}\rvert$.
Its sequential variant chooses candidates one after another, maximizing the number of newly covered voters in every step.
@@
where $H(t)=\sum_{j=1}^{t}1/j$, is the $t$-th harmonic number, with $H(0)=0$.
@@
This rule starts with a budget of $k/n$ per voter and an empty committee. Adding a candidate to the committee incurs a cost of $1$. In each step, MES selects a candidate that minimizes the maximum amount of budget any voter has to pay. A detailed description can be found in Algorithm 3. The budget of each voter is $b=k/n$.
@@
This rule iteratively detects the largest current EJR+ violation and adds the responsible candidate to the committee. Under GJCR, the addition of a candidate incurs a cost of $1$, which is uniformly split among all voters involved in the violation. For a formal definition, see Algorithm 4.
@@ Appendix B Omitted Proofs
In this appendix, we present the proofs missing from the main body of our paper.
@@ B.1 General Properties
Let $\alpha_{1}\leq\alpha_{2}$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$. First, it holds that $\alpha_{1}$-$\Phi$ $\implies\alpha_{2}$-$\Phi$ since increasing $\alpha$ can only decrease the number of cohesive groups, which therefore can only make the axiom weaker. Therefore, $\alpha_{1}$-EJR+ $\implies$ $\alpha_{2}$-EJR+. Let $\alpha_{2}$-EJR+ be satisfied by committee $W$. Assume for contradiction that $\alpha_{2}$-EJR is violated with witness $(N^{\prime},C^{\prime},\ell)$ where $N^{\prime}\subseteq N$ and $C^{\prime}\subseteq C$. For all $v\in N^{\prime}$, $\lvert A_{v}\cap W\rvert$$\ell$ and due to cohesiveness of $N^{\prime}$, there is a candidate $c\in C^{\prime}\setminus W$. Since $\lvert N_{c}\rvert\geq\alpha_{2}\cdot\ell\frac{n}{k}$, this candidate causes an $\alpha_{2}$-EJR+ violation; a contradiction.
@@
First, we show that $\alpha_{1}$-EJR+ does not imply $\alpha_{3}$-JR. For this, consider the following instance. Let $\alpha_{2}=\frac{a}{b}$, $N=\{1,\dots b\}$ and $C=\{c_{1},c_{2}\}$, $k=1$, with $A_{v}=\{c_{1}\}$ for $v\in\{1,\dots,a\}$. Then, the committee $W=\{c_{2}\}$ satisfies $\alpha_{1}$-EJR+. This follows from the fact that since
$\alpha_{1}\cdot\frac{n}{k}>\alpha_{2}\cdot\frac{n}{k}=\frac{a}{b}\cdot b=a,$
there does not exist any cohesive groups for $\alpha_{1}$ in the instance. On the other hand, the voters $\{1,\dots,a\}$ together with candidate $c_{1}$ are witnessing an $\alpha_{2}$-JR violation.
The same example can also be used to show that $\alpha_{1}$-EJR+ does not imply $\alpha_{2}$-EJR and that $\alpha_{2}$-EJR does not imply $\alpha_{3}$-JR.
Next, we show that $\alpha_{2}$-EJR does not imply $\alpha_{1}$-EJR+. So let $\alpha_{1}=\frac{a_{1}}{b_{2}}$ and $\alpha_{2}=\frac{a_{2}}{b_{2}}$ with $b_{1}\geq 2$. We create the instance $N=[(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})]$, with $C=\{c_{1},\dots,c_{2\cdot a_{1}\cdot b_{2}}\}\cup\{c\}$ and $k=2\cdot a_{1}\cdot b_{2}$ and $A_{v_{i}}=\{c,c_{i}\}$ for $i\in[2\cdot a_{1}\cdot b_{2}]$. Thus, $\alpha_{1}\cdot\frac{n}{k}=\frac{a_{1}}{b_{1}}\cdot\frac{(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})}{2\cdot a_{1}\cdot b_{2}}=(a_{1}\cdot b_{2})$ and $\alpha_{2}\cdot\frac{n}{k}=\frac{a_{2}}{b_{2}}\cdot\frac{(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})}{2\cdot a_{1}\cdot b_{2}}=a_{2}\cdot b_{1}$.
We claim that the committee $W=\{c_{1},\dots,c_{2\cdot a_{1}\cdot b_{2}}\}$ satisfies $\alpha_{2}$-EJR but violates $\alpha_{1}$-EJR+. First, observe that the committee violates $\alpha_{1}$-EJR+. This follows, since $\lvert N_{c}\rvert=2\cdot\alpha_{1}\cdot\frac{n}{k}$, no voter in $N$ approves $2$ candidates in $W$ and $c\not\in W$. On the other hand, we claim that there does not exist an $(\alpha,2)$-cohesive group. Since $\lvert A_{v}\cap A_{v^{\prime}}\rvert\leq 1$ for any $v\neq v^{\prime}\in N$ an $(\alpha,2)$-cohesive group can contain at most one voter. But as computed above, an $(\alpha,2)$-cohesive group has to be of size at least $a_{2}\cdot b_{1}\geq 2$. On the other hand, every voter who approves of any candidate approves of at least one candidate in $W$. Thus, $W$ satisfies $(\alpha,2)$-EJR.
@@
Lastly, we show that $\alpha_{2}$-JR does not imply $\alpha_{1}$-EJR for $\alpha_{1}>\alpha_{2}$. Consider the following instance $N=[4],C=\{c_{1},c_{2},c_{3}\},k=2$ with $A_{i}=\{c_{1},c_{2}\}$ for $i\in[4]$. Then, $W=\{c_{1},c_{3}\}$ satisfies $\alpha_{2}$-JR for all $\alpha_{2}>0$ since every voter approves at least one candidate in $W$. On the other hand, $W$ does not satisfy EJR for any $\alpha_{1}\leq 1$ since $N$ together with $\{c_{1},c_{2}\}$ witness an $(\alpha_{1},2)$-EJR violation. ∎
@@
Consider a generalization of the example from Figure˜3. We have a set of voters $\{1,\dots,x\}$ approving candidates $c_{1}$ to $c_{k}$ and a set of voters $\{x+1,\dots,x+y\}$ approving candidate $c_{k+1}$. Without loss of generality, let $c_{1}$ to $c_{k-1}$ be in the committee $W$. If $c_{k+1}$ is also in the committee, then the voters approving candidate $c_{k}$ are ($\alpha_{1}$, k)-cohesive for an $\alpha_{1}$-value of $\frac{x}{x+y}$. If $c_{k}$ instead of $c_{k+1}$ is in the committee, then the voters approving $c_{k+1}$ are $(\alpha_{2}$, 1)-cohesive for an $\alpha_{1}$-value of $\frac{k\cdot b}{a+b}$. So the optimal $\alpha$-value with respect to EJR is
@@
By setting $x=k\cdot y$, both values simplify to $k/(k+1)$. On the other hand, there exists a committee such that every voter approves at least one candidate, for instance, the committee $W=\{c_{k+1},c_{1},\dots,c_{k-1}\}$, the optimal $\alpha$-value is equal to $1/n$. The optimal committee with respect to $\alpha$-JR $\{c_{1},c_{k-1}\}$, resulting in an optimal $\alpha$-value of $1/n$ where $n=x+y=k\cdot y$. Therefore, the additive gap between JR and EJR for the optimal $\alpha$-value tends to $k/(k+1)$. ∎
@@
In the main body of the paper Example˜4.12 gave an example which relied on tie breaking to show that many rules can return bad committees. In the following we provide an example which does not rely on tie-breaking. The arguments work the same as for Example˜4.12, with the difference that the optimal value in this instance is $2\cdot\frac{k}{n}$ achieved by the committee $W=D$.
@@
Consider the following family of instances. Let $N=[(k+1)(k+2)]$. Furthermore, let $C=B\cup D$, with $B=\{b_{1},\dots,b_{k+1}\}$ such that $N_{b_{i}}=\{(i-1)\cdot(k+2)+1,\dots,i\cdot(k+2)\}$ and $D=\{d_{1},\dots,d_{k}\}$ such that $N_{d_{i}}=\{j\in N\colon(j\mod(k+2))=i\}$ for all $i\in\{1,\dots,k+1\}$. Then each candidate in $D$ and each candidate in $B$ is approved by $k+1$ voters.

### v2

@@ 6 Restricted Domains
Since it is computationally hard to compute $\alpha^{*}_{\Phi}(I)$ for general instances $I$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$, we look into some restricted domains in which we can compute the optimal $\alpha$-value in polynomial time. These domains are well-studied restrictions to the general multiwinner voting setting. Preferences are aligned along a single dimension, such as voters and candidates positioned on a political spectrum, geographic proximity in local elections, or a preference for candidates differing along a skill or topic dimension.
@@ 6.1 Party-List Profiles
A commonly studied restricted domain is that of party-list profiles, in which voters’ preferences depend solely on candidates’ party affiliations.
@@
[…]
In the proof, we use the d’Hondt method from the apportionment setting (d’Hondt, 1882) and show that it constructs an optimal committee with regard to $\alpha^{*}_{\text{EJR+}}(I)$.
@@ 6.2 Voter-Interval Domain
Aziz et al. (2017) showed that checking whether $W$ provides EJR for an instance $I$ is coNP-complete. However, to the best of our knowledge, there is no result on efficiently finding violations in $\ell$-cohesive groups when we consider only restricted approval profiles. More specifically, we prove the following lemma, which allows us to efficiently check for an $\alpha$-EJR violation given an instance $I$ and a committee $W$.
@@
Given an instance $I$ in the VI domain, we can verify whether a committee $W$ satisfies $\alpha$-EJR in $O(n^{3}k)$.
Furthermore, we show that we can compute the optimal value $\alpha_{\text{JR}}^{*}(I)$ for a given instance $I$ in the VI domain. This shows that, by restricting voters’ preferences, we can obtain a positive result in contrast to the hardness result shown in Theorem˜5.2 for general instances.
@@
Computing $\alpha^{*}_{\text{JR}}(I)$ in the VI domain can be done in $O(n^{2}m\log n)$.
The greedy algorithm in Algorithm˜1 constructs a subset of candidates that satisfies $\alpha$-JR for a given $I$ and $\alpha$, which is cardinal-minimal.
By checking if the size of the returned committee is at most $k$, we can verify if, for this $\alpha$-value, we obtain a witness (feasible committee) satisfying $\alpha$-JR. Due to Lemma˜4.10, the number of distinct $\alpha$-values we need to consider is upper-bounded by $n$. We can go through the $\alpha$-values using binary search to find the highest value such that Algorithm˜1 returns a committee that has more than $k$ candidates, as this shows that there does not exist a committee that can satisfy $\alpha$-JR for the $\alpha$-value given in this step.
In Algorithm˜1, we use the observation that in the VI domain, every candidate $c$ has a left-most and right-most supporting voter according to the VI order. We denote these voters by $l_{c}$ and $r_{c}$, respectively. For $i\in[n]$, the algorithm considers the set $\{v_{1},\dots v_{i}\}$ and fixes an $\alpha$-JR violation, if existent, with a candidate approved by the right-most voter. Intuitively, this ensures that the violation to the left of $v_{i}$ is resolved while covering as many voters as possible to the right, thereby postponing the next potential violation as long as possible.
@@
   $W\leftarrow W\cup\{c\}$ with $c\in A_{v_{i}}$ and
     $r_{c^{\prime}}\sqsubseteq r_{c}$ for all $c^{\prime}\in A_{v_{i}}$.
@@ 6.3 Candidate-Interval Domain
Given an instance $I$ in the CI domain, we can verify whether a committee $W$ satisfies $\alpha$-EJR in $O(m^{2}kn)$.
@@
Computing $\alpha^{*}_{\text{JR}}(I)$ in the CI domain can be done in $O(m^{2}n\log n)$.
@@
For both the CI and VI domains, we leave open the question of whether a polynomial-time algorithm exists to compute the optimal $\alpha$-value for EJR, $\alpha_{\text{EJR}}^{*}(I)$, for a given instance $I$. The same holds for EJR+. The proposed left-to-right sweeping methodology does not work when $\ell>1$.
@@ 7 Empirical Evaluation
Figure 5: Distribution of optimal $\alpha$-values, $\alpha_{\text{EJR}}^{*}$, under the IC and the EUT model, based on 6400 generated instances each.
Figure 6: Cumulative distribution of the optimal $\alpha$-value, $\alpha_{\text{EJR}}^{*}$, and the $\alpha$-value achieved by MES, seq-Phragmén, $C$, and PAV under the IC and the EUT model, based on 6400 samples.
Finally, we complement our theoretical results with an empirical analysis. We generate synthetic voting instances to evaluate the resulting values of $\alpha$ and to assess the performance of existing multi-winner voting rules. Our primary focus is on $\alpha$-EJR; additional plots for $\alpha$-JR are available in the full version. Figure˜5 reports the distribution of optimal values, $\alpha^{*}_{\text{EJR}}$, under the Impartial Culture (IC) and the Euclidean Threshold (EUT) model, two commonly used probabilistic models for generating approval profiles. For each model, we consider instances of varying sizes—ranging from small to moderate—with up to $60$ voters and $15$ candidates. Both models are widely used in approval-based multiwinner voting (Bredereck et al., 2019; Faliszewski et al., 2017; Elkind et al., 2017). Further details on the parameter configurations of them are provided in Appendix˜C.
@@
In addition to our theoretical analysis in Section˜4.2, we evaluate the performance of existing voting rules on randomly generated instances. Figure˜6 demonstrates that MES, seq-Phragmén, and PAV typically achieve committees with comparatively low $\alpha$-values with respect to EJR, indicating that these rules perform well on average. Note that MES is completed with seq-Phragmén. Our experiments, however, also uncover instances in which all three rules yield committees whose $\alpha$-values exceed the optimal value by more than a factor of four, highlighting substantial worst-case deviations even in average-case data. These cases are mostly seen in larger instances. In Appendix˜C, we provide more fine-grained information on how the voting rules perform under different instance sizes. CC does not perform very well, which is unsurprising because its only goal is to satisfy as many voters as possible and because it does not satisfy EJR for all instances. In the full version of the paper, we show that, on the contrary, CC performs best out of the considered rules for $\alpha_{\text{JR}}^{*}$.
@@ 8 Conclusion
We demonstrate that our dynamic extensions of JR and EJR can lead to much fairer outcomes for various instances. Our axioms possess a clear fairness appeal, as they seek to minimize the underrepresentation of large aligned groups, which in real-world political settings may otherwise foster resentment. However, as with many other voting rules that aim to maximize an objective, such as CC or PAV, computing the optimal $\alpha$-value is NP-hard, even for JR.
While existing voting rules can exhibit poor worst-case behavior theoretically, they perform reasonably well on our artificially generated datasets. A natural direction for future work is to evaluate these rules on real-world data, for instance, using the Polkadot dataset (Boehmer et al., 2024). When restricting the set of voting instances, we obtain several positive results regarding the efficient computation and verification of $\alpha$-values. Empirical results underscore the importance of considering adaptive quotas in multiwinner approval voting and, perhaps, in other settings. Further extensions could include generalizing our framework to weakly cohesive groups or developing broader notions of cohesiveness based on justified representation functions. Such generalizations may also prove useful in continuous settings, such as budget aggregation.
Acknowledgments
We want to thank our AAMAS 2026 reviewers for their valuable feedback. Patrick and Fabian are supported by the Deutsche Forschungsgemeinschaft under grant BR 2312/14-1.
@@ Appendix A Mentioned Voting Rules
Selects all committees that maximize the CC-score, i.e. $\arg\max_{W\subseteq C,\lvert W\rvert=k}\text{CC}(W)$. The CC-score equals the number of voters who approve at least one candidate in the committee. Formally,
$\text{CC}(W)=\lvert\{v\in N:A_{v}\cap W\neq\emptyset\}\rvert$.
Its sequential variant chooses candidates one at a time, maximizing the number of newly covered voters at each step.
@@
where $H(t)=\sum_{j=1}^{t}\frac{1}{j}$, is the $t$-th harmonic number, with $H(0)=0$.
@@
This rule starts with a budget of $k/n$ per voter and an empty committee. Adding a candidate to the committee incurs a cost of $1$. At each step, MES selects a candidate that minimizes the maximum budget any voter has to pay. A detailed description can be found in Algorithm 3. Each voter’s budget is $b=k/n$.
@@
This rule iteratively detects the largest current EJR+ violation and adds the responsible candidate to the committee. Under GJCR, adding a candidate incurs a cost of $1$, which is uniformly split among all voters involved in the violation. For a formal definition, see Algorithm 4.
@@ Appendix B Omitted Proofs
In this section, we present the proofs missing from the main body of our paper.
@@ B.1 General Properties
Let $\alpha_{1}\leq\alpha_{2}$ and $\Phi\in\{\text{JR, EJR, EJR+}\}$. First, it holds that $\alpha_{1}$-$\Phi$ $\implies\alpha_{2}$-$\Phi$ since increasing $\alpha$ can only decrease the number of cohesive groups, which therefore can only make the axiom weaker. Put differently, every $\alpha_{2}$-cohesive group is also an $\alpha_{1}$-cohesive group. Therefore, $\alpha_{1}$-EJR+ $\implies$ $\alpha_{2}$-EJR+. Let $\alpha_{2}$-EJR+ be satisfied by committee $W$. Assume for contradiction that $\alpha_{2}$-EJR is violated with witness $(N^{\prime},C^{\prime},\ell)$ where $N^{\prime}\subseteq N$ and $C^{\prime}\subseteq C$. For all $v\in N^{\prime}$, $\lvert A_{v}\cap W\rvert$$\ell$ and due to cohesiveness of $N^{\prime}$, there is a candidate $c\in C^{\prime}\setminus W$. Since $\lvert N_{c}\rvert\geq\alpha_{2}\cdot\ell\frac{n}{k}$, this candidate causes an $\alpha_{2}$-EJR+ violation; a contradiction.
@@
First, we show that $\alpha_{1}$-EJR+ does not imply $\alpha_{3}$-JR. Select $a,b\in\operatorname{\mathbb{N}}$ such that $\alpha_{1}>\frac{a}{b}\geq\alpha_{3}$ and consider the following instance. Let $\alpha_{3}=\frac{a}{b}$ for $a,b\in\operatorname{\mathbb{N}}$, $N=\{v_{1},\dots v_{b}\}$ and $C=\{c_{1},c_{2}\}$, $k=1$, with $A_{v}=\{c_{1}\}$ for $v\in\{v_{1},\dots,v_{a}\}$. Then, the committee $W=\{c_{2}\}$ satisfies $\alpha_{1}$-EJR+. This follows from the fact that since
$\alpha_{1}\cdot\frac{n}{k}>\frac{a}{b}\cdot\frac{n}{k}=\frac{a}{b}\cdot b=a,$
there does not exist any cohesive groups for $\alpha_{3}$ in the instance. On the other hand, the voters $\{v_{1},\dots,v_{a}\}$ together with candidate $c_{1}$ are witnessing an $\alpha_{3}$-JR violation.
The same construction can also be used to show that $\alpha_{1}$-EJR+ does not imply $\alpha_{2}$-EJR and that $\alpha_{2}$-EJR does not imply $\alpha_{3}$-JR.
Next, we show that $\alpha_{2}$-EJR does not imply $\alpha_{1}$-EJR+. Let $\frac{a_{1}}{b_{1}}>\alpha_{1}>\alpha_{2}>\frac{a_{2}}{b_{2}}$ for some $a_{1},a_{2},b_{1},b_{2}\in\operatorname{\mathbb{N}}$. We create the instance $N=[(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})]$, with $C=\{c_{1},\dots,c_{2\cdot a_{1}\cdot b_{2}}\}\cup\{c\}$ and $k=2\cdot a_{1}\cdot b_{2}$ and $A_{v_{i}}=\{c,c_{i}\}$ for $i\in[2\cdot a_{1}\cdot b_{2}]$. Thus, $\alpha_{1}\cdot\frac{n}{k}<\frac{a_{1}}{b_{1}}\cdot\frac{(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})}{2\cdot a_{1}\cdot b_{2}}=a_{1}\cdot b_{2}$ and $\alpha_{2}\cdot\frac{n}{k}>\frac{a_{2}}{b_{2}}\cdot\frac{(2\cdot b_{1}\cdot b_{2})\cdot(a_{1}\cdot b_{2})}{2\cdot a_{1}\cdot b_{2}}=a_{2}\cdot b_{1}$.
We claim that the committee $W=\{c_{1},\dots,c_{2\cdot a_{1}\cdot b_{2}}\}$ satisfies $\alpha_{2}$-EJR but violates $\alpha_{1}$-EJR+. First, observe that the committee violates $\alpha_{1}$-EJR+. This follows, since $\lvert N_{c}\rvert=2\cdot a_{1}\cdot b_{2}>2\cdot\alpha_{1}\cdot\frac{n}{k}$, no voter in $N$ approves $2$ candidates in $W$, and $c\not\in W$. On the other hand, we claim that there does not exist an $(\alpha_{2},2)$-cohesive group. Since $\lvert A_{v}\cap A_{v^{\prime}}\rvert\leq 1$ for any $v\neq v^{\prime}\in N$ an $(\alpha_{2},2)$-cohesive group can contain at most one voter. But as computed above, an $(\alpha_{2},2)$-cohesive group has to be of size at least $2\cdot\alpha_{2}\cdot\frac{n}{k}>2\cdot a_{2}\cdot b_{1}\geq 2$. On the other hand, every voter who approves of any candidate approves of at least one candidate in $W$. Thus, $W$ satisfies $\alpha_{2}$-EJR.
@@
Lastly, we show that $\alpha_{3}$-JR does not imply $\alpha_{2}$-EJR. Consider the following instance $N=[4],C=\{c_{1},c_{2},c_{3}\},k=2$ with $A_{v}=\{c_{1},c_{2}\}$ for $v\in[4]$. Then, $W=\{c_{1},c_{3}\}$ satisfies $\alpha_{3}$-JR for all $\alpha_{3}>0$ since every voter approves at least one candidate in $W$. On the other hand, $W$ does not satisfy $\alpha_{2}$-EJR for any $\alpha_{2}\leq 1$ since $N$ together with $\{c_{1},c_{2}\}$ witness an $(\alpha_{2},2)$-EJR violation. ∎
@@
Consider a generalization of the example from Figure˜3. We have a set of voters $\{1,\dots,x\}$ approving candidates $c_{1}$ to $c_{k}$ and a set of voters $\{x+1,\dots,x+y\}$ approving candidate $c_{k+1}$. Without loss of generality, let $c_{1}$ to $c_{k-1}$ be in the committee $W$. If $c_{k+1}$ is also in the committee, then the voters approving candidate $c_{k}$ are ($\alpha_{1}$, k)-cohesive for an $\alpha_{1}$-value of $\frac{x}{x+y}$ and, therefore, witness an $\alpha_{1}$-EJR violation. If $c_{k}$ instead of $c_{k+1}$ is in the committee, then the voters approving $c_{k+1}$ are $(\alpha_{2}$, 1)-cohesive for an $\alpha_{2}$-value of $\frac{k\cdot y}{x+y}$, and witness an $\alpha_{2}$-EJR violation. So the optimal $\alpha$-value with respect to EJR is
@@
By setting $x=k\cdot y$, both values simplify to $k/(k+1)$. On the other hand, the optimal committee with respect to $\alpha$-JR, $W=\{c_{1},c_{k-1}\}$ results in an optimal $\alpha$-value of 0 as we can cover every voter by $W$. Therefore, the additive gap between JR and EJR for the optimal $\alpha$-value is $k/(k+1)$. ∎
@@
In the main body of the paper, Example˜4.12 presents an instance $I$ that relies on tie-breaking to show that many rules can return suboptimal committees regarding $\alpha^{*}(I)$. In the following, we provide an example that does not rely on tie-breaking. The argument works the same way as the one for Example˜4.12, with the difference that the optimal value in this instance is $2\cdot\frac{k}{n}$ achieved by the committee $W=D$.
@@
Consider the following family of instances. Let $N=[(k+1)(k+2)]$. Furthermore, let $C=B\cup D$, with $B=\{b_{1},\dots,b_{k+1}\}$ such that $N_{b_{i}}=\{(i-1)\cdot(k+2)+1,\dots,i\cdot(k+2)\}$ and $D=\{d_{1},\dots,d_{k}\}$ such that $N_{d_{i}}=\{j\in N\colon(j\mod(k+1))=i\}$ for all $i\in\{1,\dots,k\}$. Then each candidate in $D$ is approved by $k+2$ voters and each candidate in $B$ is approved by $k+1$ voters. Since $N_{b_{i}}\cap N_{b_{j}}=\emptyset$ for $i\neq j$ most rules would select $k$ candidates from $B$.